# Markov chain Monte Carlo without evaluating the target: an auxiliary variable approach

### v2

@@ 3.2 General methodology
Methodologically, this approach addresses the issues of proposal design’ and ‘ratio estimation’ separately. Users can generate one minibatch to guide the proposal and use the other to estimate the target ratio. We will discuss this strategy further in Section 3.3.2. The ratio $\mathbb{P}_{\theta^{\prime}}(\omega_{1})/\mathbb{P}_{\theta_{t}}(\omega_{1})$ will also be cancelled out when the distribution of $\omega_{1}$ does not depend on $\theta$.
@@ 3.3 New algorithms
We now introduce new gradient-based minibatch algorithms within our framework (Algorithm 4). We focus on sampling from a tall dataset where $\pi(\theta\mid x)\propto\pi(\theta)p_{\theta}(x)=\pi(\theta)\prod_{i=1}^{N}\mathsf{p}_{\theta}(x_{i})$. Our goal is to incorporate gradient information in the proposal design while maintaining the correct stationary distribution and minimal overhead. Each iteration only requires a minibatch of data. The high-level idea is to introduce an auxiliary variable ($\omega_{1}$ in Algorithm 4) to estimate the full gradient at a low cost. Then, we use a second auxiliary variable $\omega_{2}$, as described in Section 2, to estimate the target ratio. These two variables may be either independent or correlated, depending on the application.
@@ 3.3.1 PoissonMH with locally balanced proposal
In this section, we will adhere to all the technical assumptions of PoissonMH (Algorithm 2).
@@
The Barker’s proposal in Zanella, (2020) is of the form $Q^{g}(\theta,\mathrm{d}\theta^{\prime})\propto g\left(\pi(\theta^{\prime}\mid x)/\pi(\theta\mid x)\right)K(\theta,\mathrm{d}\theta^{\prime}),$ where $g:\mathbb{R}^{+}\xrightarrow{}\mathbb{R}^{+}$ is a balancing function satisfying $g(t)=tg(1/t)$. Here, $K$ represents a symmetric kernel. Users can choose the balancing function, with recommended options including $g(t)=\sqrt{t}$ and $g(t)=t/(1+t)$. Experiments on various discrete distributions show that its performance is competitive compared to alternative MCMC methods.
When $\theta$ is in a continuous state space, the proposal $Q^{g}(\theta,\cdot)$ is not feasible for implementation. However, one can perform a first–order Taylor expansion of $\log(\pi(\theta^{\prime}\mid x))-\log(\pi(\theta\mid x))$ with respect to $\theta$, and utilise the so-called first-order locally balanced proposal. It has the form: $Q^{(g)}(\theta,\mathrm{d}\theta^{\prime})=\prod_{i=1}^{d}Q^{(g)}_{i}(\theta,\mathrm{d}\theta^{\prime}_{i}).$ Here $Q^{(g)}_{i}$ is a one-dimensional kernel of the form
@@
Our proposal: We retain all promises and notations used in Algorithm 2. We also use the notation $\omega_{1}:=(s_{1},s_{2},\ldots,s_{N})$ and $\mathbb{P}_{\theta}(\cdot):=\otimes_{i=1}^{N}\mathsf{Poi}\left(\frac{\lambda M_{i}}{L}+\phi_{i}(\theta;x)\right)$, as defined in Example 2. For every balancing function $g$, we define the Markov transition kernels with density $q_{\omega_{1}}^{(g)}(\theta,\theta^{\prime}):=\prod_{i=1}^{d}q_{\omega_{1},i}^{(g)}(\theta,\theta_{i}^{\prime}).$ Here $q_{\omega_{1},i}^{(g)}$ is a one-dimensional density proportional to
$g\left(e^{\partial_{\theta_{i}}\log\left(\pi(\theta\mid x)\mathbb{P}_{\theta}(\omega_{1})\right)(\theta^{\prime}_{i}-\theta_{i})}\right)\mu_{i}(\theta^{\prime}_{i}-\theta_{i}),$ (3)
where $\mu_{i}$ is a symmetric density on $\mathbb{R}$. Our algorithm is summarised in Algorithm 5.
@@
1:  Initialise: Initial state $\theta_{0}$; balancing function $g$; number of iterations $T$
@@
Firstly, the proposed method is an auxiliary-variable-based variant of the locally-balanced proposal. Our proposal distribution $q_{\omega_{1}}^{(g)}$ closely resembles the first-order locally balanced proposal, with one key modification. The term $\partial_{\theta_{i}}\log\pi(\theta\mid x)$ in (2) is substituted by $\partial_{\theta_{i}}\log\left(\pi(\theta\mid x)\mathbb{P}_{\theta}(\omega_{1})\right)$ in (3). Meanwhile, standard calculation shows the proxy function $\pi(\theta\mid x)\cdot\mathbb{P}_{\theta}(\omega_{1})$ only depends on the minibatch $S$. Consequently, the cost of computing the gradient is at the same level as calculating the acceptance ratio in PoissonMH, thus not significantly increasing the cost per step. Detailed calculations are in supplementary B.
@@
Lastly, following Livingstone and Zanella, (2022), we choose $g(t)=t/(1+t)$ and $g(t)=\sqrt{t}$ in our actual implementation. We call them Poisson–Barker and Poisson–MALA respectively. These are the minibatch versions (covering both proposal generation and target ratio evaluation) of the Barker proposal and MALA. Similar to Livingstone and Zanella, (2022), these choices of $g$ allow efficient implementation, with details available in supplementary B. Their numerical performances and comparisons are examined in Section 5.
@@ 3.3.2 TunaMH with SGLD proposal
Therefore, we adopt another cost-effective strategy for utilising the gradient information, inspired by the SGLD algorithm in Welling and Teh, (2011). At each step, we first select a minibatch $B$ of $K$ data points uniformly from the entire dataset. Thus $(N/K)\sum_{i\in B}\nabla_{\theta}U_{i}(\theta_{t};x)$ is a natural estimator of the gradient of the log-posterior: $\nabla_{\theta}\log\pi(\theta\mid x)=-\nabla_{\theta}\sum_{i=1}^{N}U_{i}(\theta;x)$. Calculating this estimator has a cost that scales linearly with the minibatch size $K$, rather than with the total dataset size $N$. In the examples we considered, data size $N$ is no less than $10,000$ while $K$ is often chosen as $20$. Therefore, computing the estimated gradient is significantly cheaper compared to calculating the full gradient. Setting $\omega_{1}:=B\sim\mathsf{Unif}\{\{1,2,\ldots,N\},K\}$, our proposal has the form $q_{\omega_{1}}(\theta,\cdot)\sim\mathbb{N}\left(\theta-\frac{\epsilon^{2}}{2}\frac{N}{K}\sum_{i\in B}\nabla_{\theta}U_{i}(\theta;x),\epsilon^{2}\mathbb{I}\right),$ where $\epsilon$ is a tuning parameter. Next, we use the same strategy as TunaMH to select an independent minibatch to estimate the target ratio. Using the notations $\omega_{2}:=(s_{1},s_{2},\ldots,s_{N})$ and $\mathbb{P}_{\theta,\theta^{\prime}}(\cdot):=\otimes_{i=1}^{N}\mathsf{Poi}\left(\frac{\lambda c_{i}}{C}+\phi_{i}(\theta,\theta^{\prime};x)\right)$, our algorithm is summarised in Algorithm 6.
Algorithm 6 corresponds to Case 3 of our meta-algorithm 4. Thus, its validity directly follows from Proposition 2. Our first auxiliary variable $\omega_{1}$ is independent of the second. Therefore, the distribution $\mathbb{P}_{\theta,\theta^{\prime}}(\cdot\mid\omega_{1})$ equals $\mathbb{P}_{\theta,\theta^{\prime}}(\cdot)$. Additionally, the first variable $\omega_{1}=B$ follows a fixed distribution that does not depend on $\theta$. Consequently, the joint distribution $\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{1},\omega_{2})$ equals $\mathbb{P}(\omega_{1})\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2})$. As a result, Step 6 of Algorithm 4 simplifies to
@@
For readers already familiar with SGLD, another natural perspective is to view Algorithm 6 as a minibatch Metropolised version of SGLD. If each iteration of Algorithm 6 skips Steps 5–7 for the acceptance–rejection correction, it becomes exactly the SGLD algorithm described in Welling and Teh, (2011) with a fixed step size. The SGLD algorithm is a noisier but much faster version of the Unadjusted Langevin Algorithm (ULA). Both ULA and SGLD with a fixed step size have systematic bias, as they do not converge to the target distribution $\pi(\theta\mid x)$ even after an infinite number of steps (Vollmer et al.,, 2016). Introducing a straightforward Metropolis–Hastings acceptance–rejection step can ensure the correct stationary distribution but requires full-dataset evaluation, which undermines the advantage of SGLD. Therefore, deriving an acceptance–rejection version of SGLD using minibatch data is considered an open problem. To quote from Welling and Teh, (2011): Interesting directions of future research includes deriving a MH rejection step based on mini-batch data $\ldots$
Algorithm 6 offers a solution to this problem under the same assumptions as Algorithm 3. It eliminates the systematic bias in SGLD by incorporating an acceptance–rejection step based on the approach used in TunaMH. Relaxing the technical assumptions or generalizing to posterior distributions with non-i.i.d. model likelihoods are intriguing open questions.
Algorithm 6 TunaMH–SGLD
1:  Initialise: Initial state $\theta_{0}$; batch size $K$; step size $\epsilon$; number of iterations $T$
@@ 4 Theory
This section presents our theoretical analysis. We will begin by examining the meta-algorithm 4 and then apply our findings to specific algorithms. Proposition 2 has already shown the reversibility of the proposed method. Therefore, this section will focus on understanding the convergence speed. Our agenda is to compare the Markov chain in Algorithm 4, with transition kernel denoted by $\mathbb{P}_{\mathsf{aux}}$ with two relevant chains $\mathbb{P}_{\mathsf{ideal}}$ and $\mathbb{P}_{\mathsf{MwG}}$. These two chains need not be implementable, but are easier to study as a stochastic process.
Fix $\theta\in\Theta$, define the idealised proposal $Q_{\mathsf{ideal}}(\theta,\mathrm{d}\theta^{\prime}):=\mathbb{E}_{\omega_{1}\sim\mathbb{P}_{\theta}}[Q_{\omega_{1}}(\theta,\mathrm{d}\theta^{\prime})]$, with density $q_{\mathsf{ideal}}(\theta,\theta^{\prime})=\int_{\Omega_{1}}q_{\omega_{1}}(\theta,\theta^{\prime})\mathbb{P}_{\theta}(\omega_{1})\lambda_{1}(\mathrm{d}\omega_{1})$. The idealised chain $\mathbb{P}_{\mathsf{ideal}}$ refers to the standard Metropolis-Hastings algorithm with proposal $Q_{\mathsf{ideal}}$. Detailed description of $\mathbb{P}_{\mathsf{ideal}}$ is given in Algorithm 7 in supplementary A.4. The second chain, $\mathbb{P}_{\mathsf{MwG}}$, is defined as the transition kernel for Case 2, i.e., $\omega_{1}=\omega_{2}$ in Algorithm 4. A crucial feature is that $\mathbb{P}_{\mathsf{MwG}}$ can equivalently be viewed as a Metropolis–within–Gibbs algorithm. It is the marginal chain for the $\theta$ component, targeting an augmented distribution $\pi(\theta,\omega_{1}\mid x):=\pi(\theta\mid x)\mathbb{P}_{\theta}(\omega_{1})$. At each step, users first generate $\omega_{1}$ from $\pi(\cdot\mid\theta,x)=\mathbb{P}_{\theta}(\cdot)$. They then use the proposal $q_{\omega_{1}}(\theta,\cdot)$ to draw $\theta^{\prime}$, targeting $\pi(\theta^{\prime}\mid\omega_{1},x)\propto\pi(\theta^{\prime}\mid x)\mathbb{P}_{\theta}(\omega_{1})$, and implement an acceptance–rejection step. This perspective allows us to apply recent results, such as Qin et al., (2023), in our analysis.
@@ 4.1 Peskun’s ordering
The following result generalises existing results that use only one auxiliary variable, such as Lemma 1 in Wang, (2022), Section 2.3 in Nicholls et al., (2012), and Section 2.1 in Andrieu et al., 2018a . Nevertheless, the proof idea is very similar and is given in supplementary 4.1.
@@
This lemma shows that $\mathbb{P}_{\mathsf{aux}}$ is always less likely to move compared to $\mathbb{P}_{\mathsf{ideal}}$ and $\mathbb{P}_{\mathsf{MwG}}$. Lemma 1 implies $\mathbb{P}_{\mathsf{aux}}\prec\mathbb{P}_{\mathsf{MwG}}\prec\mathbb{P}_{\mathsf{ideal}}$ according to Peskun’s ordering (Peskun,, 1973). This means $\mathbb{P}_{\mathsf{MwG}}-\mathbb{P}_{\mathsf{aux}}$ and $\mathbb{P}_{\mathsf{ideal}}-\mathbb{P}_{\mathsf{MwG}}$ are non-negative operators, and the asymptotic variance for any $f$ under $\mathbb{P}_{\mathsf{ideal}}$ is no greater than under $\mathbb{P}_{\mathsf{aux}}$. For more results, see Tierney, (1998).

### v3

@@ 3.2 General methodology
This approach addresses the issues of proposal design’ and ‘ratio estimation’ separately. Users can generate one minibatch to guide the proposal and use the other to estimate the target ratio. We will discuss this strategy further in Section 3.3.2. The ratio $\mathbb{P}_{\theta^{\prime}}(\omega_{1})/\mathbb{P}_{\theta_{t}}(\omega_{1})$ will also be cancelled out when the distribution of $\omega_{1}$ does not depend on $\theta$.
@@ 3.3 New algorithms
We now introduce new gradient-based minibatch MCMC algorithms within our framework (Algorithm 4). We focus on sampling from a tall dataset where $\pi(\theta\mid x)\propto\pi(\theta)p_{\theta}(x)=\pi(\theta)\prod_{i=1}^{N}\mathsf{p}_{\theta}(x_{i})$. Our goal is to incorporate gradient information in the proposal design while maintaining the correct stationary distribution and minimal overhead. Each iteration only requires a minibatch of data. The high-level idea is to introduce an auxiliary variable ($\omega_{1}$ in Algorithm 4) to estimate the full gradient at a low cost. Then, we use a second auxiliary variable $\omega_{2}$, as described in Section 2, to estimate the target ratio.
@@ 3.3.1 PoissonMH with locally balanced proposal
Here, we will adhere to all the technical assumptions of PoissonMH (Algorithm 2).
@@
The Barker’s proposal is of the form $Q^{g}(\theta,\mathrm{d}\theta^{\prime})\propto g\left(\pi(\theta^{\prime}\mid x)/\pi(\theta\mid x)\right)K(\theta,\mathrm{d}\theta^{\prime}),$ where $g:\mathbb{R}^{+}\xrightarrow{}\mathbb{R}^{+}$ is a balancing function satisfying $g(t)=tg(1/t)$. Here, $K$ represents a symmetric kernel. Users can choose the balancing function, with recommended options including $g(t)=\sqrt{t}$ and $g(t)=t/(1+t)$. Experiments on various discrete distributions show that its performance is competitive compared to alternative MCMC methods.
When $\theta$ is in a continuous state space, the proposal $Q^{g}(\theta,\cdot)$ is not feasible for implementation. However, one can perform a first–order Taylor expansion of $\log(\pi(\theta^{\prime}\mid x))-\log(\pi(\theta\mid x))$ with respect to $\theta$, and utilize the so-called first-order locally balanced proposal. It has the form: $Q^{(g)}(\theta,\mathrm{d}\theta^{\prime})=\prod_{i=1}^{d}Q^{(g)}_{i}(\theta,\mathrm{d}\theta^{\prime}_{i}).$ Here $Q^{(g)}_{i}$ is a one-dimensional kernel of the form
@@
Our proposal: We retain all promises and notations used in Algorithm 2. We also use the notation $\omega_{1}:=(s_{1},s_{2},\ldots,s_{N})$ and $\mathbb{P}_{\theta}(\cdot):=\otimes_{i=1}^{N}\mathsf{Poi}\left(\frac{\lambda M_{i}}{L}+\phi_{i}(\theta;x)\right)$, as defined in Example 2. For every balancing function $g$, we define the Markov transition kernels with density $q_{\omega_{1}}^{(g)}(\theta,\theta^{\prime}):=\prod_{i=1}^{d}q_{\omega_{1},i}^{(g)}(\theta,\theta_{i}^{\prime}).$ Here $q_{\omega_{1},i}^{(g)}$ is a one-dimensional density proportional to $g\left(e^{\partial_{\theta_{i}}\log\left(\pi(\theta\mid x)\mathbb{P}_{\theta}(\omega_{1})\right)(\theta^{\prime}_{i}-\theta_{i})}\right)\mu_{i}(\theta^{\prime}_{i}-\theta_{i}),$ where $\mu_{i}$ is a symmetric density on $\mathbb{R}$. Our algorithm is summarized in Algorithm 5. Step 3 uses the Poisson minibatch sampling method from Zhang and De Sa, (2019), enabling faster implementation than sampling each data point independently. See Supplementary B for implementation details.
@@
1:  Initialize: Initial state $\theta_{0}$; balancing function $g$; number of iterations $T$
@@
Firstly, the proposed method is an auxiliary-variable-based variant of the locally-balanced proposal. Our proposal distribution $q_{\omega_{1}}^{(g)}$ closely resembles the first-order locally balanced proposal, with one key modification. The term $\partial_{\theta_{i}}\log\pi(\theta\mid x)$ in (2) is substituted by $\partial_{\theta_{i}}\log\left(\pi(\theta\mid x)\mathbb{P}_{\theta}(\omega_{1})\right)$. Meanwhile, standard calculation (provided in supplementary C) shows the proxy function $\pi(\theta\mid x)\cdot\mathbb{P}_{\theta}(\omega_{1})$ only depends on the minibatch $S$. Consequently, the cost of computing the gradient is at the same level as calculating the acceptance ratio in PoissonMH, thus not significantly increasing the cost per step.
@@
Lastly, following Livingstone and Zanella, (2022), we choose $g(t)=t/(1+t)$ and $g(t)=\sqrt{t}$ in our actual implementation. We call them Poisson–Barker and Poisson–MALA respectively. These are the minibatch versions (covering both proposal generation and target ratio evaluation) of the Barker proposal and MALA. These choices of $g$ allow efficient implementation, with details available in supplementary C. Their numerical performances and comparisons are examined in Section 5.
@@ 3.3.2 TunaMH with SGLD proposal
Therefore, we adopt another cost-effective strategy for utilising the gradient information, inspired by the SGLD algorithm in Welling and Teh, (2011). At each step, we first select a minibatch $B$ of $K$ data points uniformly from the entire dataset. Thus $(N/K)\sum_{i\in B}\nabla_{\theta}U_{i}(\theta_{t};x)$ is a natural estimator of the gradient of the log-posterior: $\nabla_{\theta}\log\pi(\theta\mid x)=-\nabla_{\theta}\sum_{i=1}^{N}U_{i}(\theta;x)$. Calculating this estimator has a cost that scales linearly with the minibatch size $K$, rather than with the total dataset size $N$. In the examples we considered, data size $N$ is no less than $10,000$ while $K$ is often chosen as $20$. Therefore, computing the estimated gradient is significantly cheaper compared to calculating the full gradient. Setting $\omega_{1}:=B\sim\mathsf{Unif}\{\{1,2,\ldots,N\},K\}$, our proposal has the form $q_{\omega_{1}}(\theta,\cdot)\sim\mathbb{N}\left(\theta-\frac{\epsilon^{2}}{2}\frac{N}{K}\sum_{i\in B}\nabla_{\theta}U_{i}(\theta;x),\epsilon^{2}\mathbb{I}\right),$ where $\epsilon$ is a tuning parameter. Next, we use the same strategy as TunaMH to select an independent minibatch to estimate the target ratio. Using the notations $\omega_{2}:=(s_{1},s_{2},\ldots,s_{N})$ and $\mathbb{P}_{\theta,\theta^{\prime}}(\cdot):=\otimes_{i=1}^{N}\mathsf{Poi}\left(\frac{\lambda c_{i}}{C}+\phi_{i}(\theta,\theta^{\prime};x)\right)$, our algorithm is summarized in Algorithm 6. Step 4 uses the Poisson minibatch sampling method from Zhang et al., (2020), enabling faster implementation than sampling each data point independently. See Supplementary B for implementation details.
Algorithm 6 corresponds to Case 3 of our meta-algorithm 4. Thus, its validity directly follows from Proposition 2. Our first auxiliary variable $\omega_{1}$ is independent of the second. Therefore, the distribution $\mathbb{P}_{\theta,\theta^{\prime}}(\cdot\mid\omega_{1})=\mathbb{P}_{\theta,\theta^{\prime}}(\cdot)$. Additionally, the first variable $\omega_{1}=B$ follows a fixed distribution that does not depend on $\theta$. Consequently, the joint distribution $\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{1},\omega_{2})$ equals $\mathbb{P}(\omega_{1})\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2})$. As a result, Step 6 of Algorithm 4 simplifies to
@@
For readers already familiar with SGLD, another natural perspective is to view Algorithm 6 as a minibatch Metropolized version of SGLD. If each iteration of Algorithm 6 skips Steps 5–7 for the acceptance–rejection correction, it becomes exactly the SGLD algorithm described in Welling and Teh, (2011) with a fixed step size. The SGLD algorithm is a noisier but much faster version of the Unadjusted Langevin Algorithm (ULA). Both ULA and SGLD with a fixed step size have systematic bias, as they do not converge to the target distribution $\pi(\theta\mid x)$ even after an infinite number of steps (Vollmer et al.,, 2016). Introducing a straightforward Metropolis–Hastings acceptance–rejection step can ensure the correct stationary distribution but requires full-dataset evaluation, which undermines the advantage of SGLD. Therefore, deriving an acceptance–rejection version of SGLD using minibatch data is considered an open problem. To quote from Welling and Teh, (2011): “Interesting directions of future research includes deriving a MH rejection step based on mini-batch data $\ldots$ Algorithm 6 offers a solution to this problem under the same assumptions as Algorithm 3.
Algorithm 6 Tuna–SGLD
1:  Initialize: Initial state $\theta_{0}$; batch size $K$; step size $\epsilon$; number of iterations $T$
@@ 4 Theory
This section presents our theoretical analysis. We will begin by studying the meta-algorithm 4 and then apply our results to specific algorithms. Since Proposition 2 has already shown the reversibility, this section will focus on understanding the convergence speed. Our agenda is to compare the Markov chain in Algorithm 4, with transition kernel denoted by $\mathbb{P}_{\mathsf{aux}}$ with two relevant chains $\mathbb{P}_{\mathsf{ideal}}$ and $\mathbb{P}_{\mathsf{MwG}}$. These two chains need not be implementable, but are easier to study as a stochastic process.
Fix $\theta\in\Theta$, define the idealized proposal $Q_{\mathsf{ideal}}(\theta,\mathrm{d}\theta^{\prime}):=\mathbb{E}_{\omega_{1}\sim\mathbb{P}_{\theta}}[Q_{\omega_{1}}(\theta,\mathrm{d}\theta^{\prime})]$, with density $q_{\mathsf{ideal}}(\theta,\theta^{\prime})=\int_{\Omega_{1}}q_{\omega_{1}}(\theta,\theta^{\prime})\mathbb{P}_{\theta}(\omega_{1})\lambda_{1}(\mathrm{d}\omega_{1})$. The idealized chain $\mathbb{P}_{\mathsf{ideal}}$ is the standard Metropolis-Hastings algorithm with proposal $Q_{\mathsf{ideal}}$. Detailed description of $\mathbb{P}_{\mathsf{ideal}}$ is given in Algorithm 7 in supplementary A.4. The second chain, $\mathbb{P}_{\mathsf{MwG}}$, is defined as the transition kernel for Case 2, i.e., $\omega_{1}=\omega_{2}$ in Algorithm 4. A crucial feature is that $\mathbb{P}_{\mathsf{MwG}}$ can equivalently be viewed as a Metropolis–within–Gibbs algorithm. It is the marginal chain for the $\theta$ component, targeting an augmented distribution $\pi(\theta,\omega_{1}\mid x):=\pi(\theta\mid x)\mathbb{P}_{\theta}(\omega_{1})$. At each step, users first generate $\omega_{1}$ from $\pi(\cdot\mid\theta,x)=\mathbb{P}_{\theta}(\cdot)$. They then use the proposal $q_{\omega_{1}}(\theta,\cdot)$ to draw $\theta^{\prime}$, targeting $\pi(\theta^{\prime}\mid\omega_{1},x)\propto\pi(\theta^{\prime}\mid x)\mathbb{P}_{\theta}(\omega_{1})$, and implement an acceptance–rejection step. This perspective allows us to apply recent results Qin et al., (2023) in our analysis.
@@ 4.1 Peskun’s ordering
The following result generalizes existing results that use only one auxiliary variable, such as Lemma 1 in Wang, (2022), Section 2.3 in Nicholls et al., (2012), and Section 2.1 in Andrieu et al., 2018a . The proof is very similar and is given in supplementary 4.1.
@@
This lemma shows that $\mathbb{P}_{\mathsf{aux}}$ is always less likely to move compared to $\mathbb{P}_{\mathsf{ideal}}$ and $\mathbb{P}_{\mathsf{MwG}}$. Lemma 1 implies $\mathbb{P}_{\mathsf{aux}}\prec\mathbb{P}_{\mathsf{MwG}}\prec\mathbb{P}_{\mathsf{ideal}}$ according to Peskun’s ordering (Peskun,, 1973). This implies that the asymptotic variance for any $f$ under $\mathbb{P}_{\mathsf{ideal}}$ is bounded above by that under $\mathbb{P}_{\mathsf{MwG}}$ and $\mathbb{P}_{\mathsf{aux}}$. For further implications, see Tierney, (1998).